Determine the possible values and construct the probability distribution for the random variables described in each of the following situations. Draw the corresponding histogram for aech probabilty


2. A shipment of five computers contain two that are defective. If a reatailer receives three of these compiters at random, list the elements of the sample space S using the letters D and N for defective and non-defective computers, respectively. To each sample point assign a value x of the random variable x representing the number of computers purchased by the retailer which are defective.


1
Expert's answer
2021-05-07T09:51:28-0400

Solution:

Given, N denotes the number of non-defective computers and D denotes defective.


"\begin{array}{l}
\text { 3N, 0D: } \\
\begin{array}{l}
\text { X=0 } \\
\text { P }(3 \mathrm{~N}, 0 \mathrm{D})
\end{array}
\end{array}"

"=^3C_3\times(^2C_0/ ^5C_3)=1\times(1/10)=0.1"


"2 \mathrm{~N}, 1 \mathrm{D}:\\
\mathrm{X}=1\\
\mathbf{P}(2 \mathbf{N}, \mathbf{1} \mathbf{D})"

"=^3C_2\times(^2C_1/ ^5C_3)=3\times(2/10)=0.6"


"1 \mathrm{~N}, 2 \mathrm{D}:\\
\begin{array}{l}\\
\mathrm{X}=2 \\
\qquad \begin{array}{r}
 \\

\end{array}
\end{array}"

"\mathrm{P}(\mathbf{1 N}, 2 \mathrm{D})=^3C_1\times(^2C_2/ ^5C_3)=3\times(1/10)=0.3"

Histogram: